The lines 3x –4y + 4 = 0 and 6x –8y –7 = 0 are tangents to the same circle. The radius of the circle is____________________________________________

Grade:11

1 Answers

Anshuman Chaurasia
6 years ago
On taking 2 common from the equation of second line it is clearly visible that both the tangents are parallel to each other, thus we can safely say that the perpendicular distance between the two lines is equal to the diameter of the circle.Thus, using the formula of distance b/w two parallel lines,D= |c1-c2|/√(a^2+b^2)On applying values,D=3/2 So radius will be R=3/4